Real World Financlals The table below contalns selected Information from recent financlal statements of The Home Depot, Inc., and Lowes Companles, Inc., two companles In the home Improvement retall Industry (S In millions) Home Depot Lowes Net sales Cost of goods sold Year-end inventory 1/31/16 2/1/15 1/29/16 1/30/15 $79,122 $75,964 $53,727 $5e,831 51,73248,92 29,934 33,264 1e,94 1e,59e ,157 8,63e Industry Averages: Gross profit ratio Inventory turnover ratio Average days in inventory 33 % 3.9 times 94 days Required Calculate the gross profit ratio, the Inventory turnover ratio, and the average days In Inventory for the two companies for thelr fiscal years ending In 2016. (Use 365 days a year. Round Inventory turnover ratlo to 2 declmal places. Enter dollar amounts In milllions. Round the doller answer values to the neerest whole dollers.) oss Profit Ratio Choose Numerator: Choose Denominator Gross Profit Ratio Gross profit ratio HOME DEPOT LOWES Inventory Turnover Ratio Choose Numerator: Choose Denominator Invent Turnover Ratio Inventory turnover ratio HOME DEPOT LOWES Average Days in Inventory Choose Numerator: Choose Denominator -Average Days in Invent Average days in inventory HOME DEPOT LOWES
Answer #1

Gross Profit ratio

Gross profit = (Sales - COGS) / Sales = Gross profit ratio
Home depot (79122 - 51732) / 79122 = 34.62%
Lowe's (53727 - 29934) / 53727 = 44.28%

Inventory Turnover ratio

COGS / Average Inventory = Inventory Turnover Ratio
Home depot 51732 / (10940+10500) / 2 = 4.83 times
Lowe's 29934 / (9157+8630) / 2 = 3.37 times

Average Days in Inventory

365 / Inventory Turnover = Average Days in Inventory
Home depot 365 / 4.83 = 75.64 days
Lowe's 365 / 3.37 = 108.44 days